OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality y 2x + 4 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (swissgirl):

ok but what is the inequality is it < or >

OpenStudy (swissgirl):

is it \( y \leq 2x+4 \) or \( y \geq sx+4 \)?

OpenStudy (anonymous):

I'm just guessing, nvm. thanks though! (:

OpenStudy (swissgirl):

its kinda easy to solve if u gave us a proper equaltion

OpenStudy (anonymous):

idk thats what it said..

OpenStudy (swissgirl):

so it just said y 2x+4 ?

jimthompson5910 (jim_thompson5910):

Most likely it's either \[\Large y \le 2x+4\] or \[\Large y \ge 2x+4\] but the symbol didn't transfer

jimthompson5910 (jim_thompson5910):

If its \[\Large y \le 2x+4\] , you shade below the solid line If it's \[\Large y \ge 2x+4\] , you shade above the solid line